Alex is twenty years older than Maria. Maria is three years old. How many years form now will he be three times as old as her?

Answer by TimothyLamb(4379)

---
x = Alex
y = years from now when Alex will be 3 times as old as Maria
---
x = 20 + 3
x = 23
---
23 + y = 3(3 + y)
23 + y = 9 + 3y
2y = 14
y = 7
---
answer:
7 years from now Alex will be 3 times as old as Maria
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php